PROGRESS OF THE JILA ELECTRON EDM EXPERIMENT

HUANQIAN LOH, KEVIN C. COSSEL, MATT GRAU, \underline{DANIEL N. GRESH}, KANG-KUEN NI, JUN YE, and ERIC A. CORNELL, JILA, National Institute of Standards and Technology and University of Colorado Department of Physics, University of Colorado, Boulder, Colorado 80309-0440, USA.

Molecules can be advantageous for the search for the electron electric dipole moment (eEDM) due to the large effective electric field experienced by a bound, unpaired electron. Furthermore, the closely-spaced states of opposite parity make the molecules easy to polarize in the lab frame. The JILA eEDM experiment currently uses HfF^{+} molecules in an ion trap to achieve long coherence times to reduce systematics.^{a} When an electric field is applied the eEDM signal is proportional to the shift in energy splitting between two Zeeman levels in a low-lying, metastable ^{3}Δ_{1} state. We have previously shown efficient preparation of trapped HfF^{+} molecules in the rovibronic ground state, X^{1}Σ^{+}(v = 0, J = 0).^{b} Here, we demonstrate coherent transfer of population from the ground state to the a^{3}Δ_{1}(v = 0, J = 1) state through an intermediate ^{3}Π_{0+} state and efficient state read-out using photodissociation. In addition, we have begun to take spectroscopy data of the hyperfine and Zeeman structure of the eEDM science state in the presence of a rotating bias electric field and a magnetic field.

aA. E. Leanhardt et. al., Journal of Molecular Spectroscopy 270, 1-25 (2011).

bH. Loh et. al., Journal of Chemical Physics 135, 154308 (2011).